OpenStudy (anonymous):

1.7k-2(1.5k-2)=5(.26k+0.8)???

OpenStudy (saifoo.khan):

k=0

OpenStudy (anonymous):

can you show the work plz? :(

OpenStudy (saifoo.khan):

Sure, just a sec

OpenStudy (anonymous):

thanks :)

OpenStudy (saifoo.khan):

1.7k - 3k + 4 = 1.3k + 4 Can u solve further?

OpenStudy (anonymous):

you got me confused.... how did you get 3k, 4 and 1.3k?

OpenStudy (saifoo.khan):

just hold on a sec, let me show u more detail.

OpenStudy (anonymous):

lol okay

OpenStudy (saifoo.khan):

getting it now?

OpenStudy (anonymous):

yes, and after thiss? 1.7k - 3k + 4 = 1.3k + 4 do i have to combine like terms?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (saifoo.khan):

please hit good answer.

OpenStudy (anonymous):

already did :) so i subract 1.7k-3k-1.3k?

OpenStudy (saifoo.khan):

1.7k-3k-1.3k = 0 subtract them, u will get the answer zero

OpenStudy (anonymous):

i didnt get zero!!!???

OpenStudy (saifoo.khan):

c'mon!

OpenStudy (saifoo.khan):

1.7k-3k-1.3k = 0 1.7k - 4.3k = 0 -2.6k = 0 k = 0/-2.6 k = 0

OpenStudy (anonymous):

OHHHHHHHHHHH!!!!

OpenStudy (anonymous):

lmfao im sorry,,,,,

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

ur very smart!

OpenStudy (saifoo.khan):

Thankyou, Thankyou! ^_^

OpenStudy (anonymous):

ur welcome >:D